# Performance evaluation of SLAM-ASR: The Good, the Bad, the Ugly, and the Way Forward

### v1

@@ V Conclusions
In this paper, we address the question of what is and what is not advisable when working with a recent, widely adopted LLM-based ASR solution, i.e., SLAM-ASR. By ablation, we were able to identify the good, the bad, and the ugly aspects of this ASR paradigm in which both the speech foundation model and the LLM are frozen, and connected through a trainable linear projector.
The Good: SLAM-ASR is efficient, both in computation and data usage, while delivering competitive in-domain performance. Its flexibility allows easy integration of various speech encoders and LLMs, the only trainable component being a small linear projector, a great advantage when computational resources are scarce.
The Bad: Like many other LLM-based ASR systems, SLAM-ASR tends to strongly overfit the training domain. This is a clear disadvantage compared to traditional ASR models, which have better generalization capabilities and are less sensitive to domain shifts and speech perturbation. Our experiments show that SLAM-ASR can easily go off the rails and hallucinate (e.g., repetitions or unrelated text) when simple perturbations are applied to the speech signal.
The Ugly: Unlike LLM-based ASR approaches that use LoRA for fine-tuning the LLM, the SLAM-ASR architecture lacks clear evidence that the projector is learning an alignment between speech and text, rather than some other spurious correlation. Our analysis shows that the alignment is less evident without LoRA during training, resulting in gibberish output when speech tokens are mapped to text tokens.

### v2

@@ V Conclusions
We have investigated what is and what is not advisable when working with a recent, widely adopted LLM-based ASR solution, i.e., SLAM-ASR. By ablation, we were able to identify the good, the bad, and the ugly aspects of this ASR paradigm in which both the speech foundation model and the LLM are frozen, and connected through a trainable projector.
The Good: SLAM-ASR is efficient, both in computation and data usage, while delivering competitive in-domain performance. Its flexibility allows easy integration of various speech encoders and LLMs, the only trainable component being a simple projector, a great advantage when computational resources or data are scarce.
The Bad: Like many other LLM-based ASR systems, SLAM-ASR tends to strongly overfit the training domain. This is a clear disadvantage compared to traditional ASR models, which have better generalization capabilities and are less sensitive to domain shifts and speech perturbation. Our experiments show that SLAM-ASR can easily go off the rails and hallucinate words (e.g., repetitions or unrelated text) when simple perturbations are applied to the speech signal.
The Ugly: Unlike LLM-based ASR approaches that use LoRA for fine-tuning the LLM, the SLAM-ASR architecture lacks clear evidence that the projector is learning an alignment between speech and text, rather than some other spurious correlation. Our analysis shows that the alignment is less evident without LoRA, resulting in gibberish output when speech tokens are mapped to text tokens.